Differentiate the function 𝑓(π‘₯) = 5π‘₯ 2 + 2 using the first principle


1
Expert's answer
2021-11-22T05:43:20-0500
"f'(x)=\lim\limits_{h\to 0}\dfrac{f(x+h)-f(x)}{h}"

"=\lim\limits_{h\to 0}\dfrac{5(x+h)^2+2-(5x^2+2)}{h}"

"=\lim\limits_{h\to 0}\dfrac{5x^2+10xh+5h^2+2-5x^2-2}{h}"

"=\lim\limits_{h\to 0}(10x+5h)=10x+0=10x"

"f'(x)=(5x^2+2)' =10x"